### v1

@@ 5.3.2 Baseline settings for the portfolio optimization problem
We use DowJones dataset in Bruni et al. (2016) with $K_{\text{a}}=24$. The true return $\bm{r}^{(i)}_{\textrm{true}}$ at time period $i$ is a vector with each element equals the true return value corresponding to the dataset. The predicted return of asset $k$ from source $h$ at time period $i$ is sampled from a truncated normal distribution with mean $(r^{(i)}_{\textrm{true},k}+\mu_{h,k})$ and variance $\sigma^{2}_{h,k}$ within interval $(-1,1)$, which means no asset can lose more than $100\%$ of its value. We set $\alpha=20\%$ and $\rho=10$ in all numerical experiments. The step size for min-max error trust update is $\Delta t=0.01$. The update rate is $\eta=10^{2}$ for exponential error trust update as well as variable-share error trust update, and the share portion is $\beta=0.5$ in variable-share trust update. We conduct $30$ trials with distinct random seeds for each trust update algorithm to observe common patterns during the trust update process.
@@ 5.3.3 Baseline results for the portfolio optimization problem
MR-DRO(Min-max) $0.018\pm 8.223\times 10^{-6}$ $0.002\pm 0.233\times 10^{-3}$ $472.177\pm 5.608$
MR-DRO(Exponential) $0.018\pm 8.225\times 10^{-6}$ $0.002\pm 0.218\times 10^{-3}$ $444.359\pm 7.746$
MR-DRO(Variable-share) $0.018\pm 8.224\times 10^{-6}$ $0.002\pm 0.219\times 10^{-3}$ $489.041\pm 6.029$
DRO (h1) $0.018\pm 1.324\times 10^{-5}$ $0.004\pm 0.703\times 10^{-3}$ $247.786\pm 3.377$
DRO (h2) $0.018\pm 8.312\times 10^{-6}$ $0.004\pm 0.799\times 10^{-3}$ $251.808\pm 3.679$
DRO (h3) $0.018\pm 1.001\times 10^{-5}$ $0.005\pm 0.673\times 10^{-3}$ $248.473\pm 3.262$
DRO (h4) $0.018\pm 1.071\times 10^{-5}$ $0.004\pm 0.635\times 10^{-3}$ $248.997\pm 3.396$
@@
We also performed an out-of-sample performance analysis. The out-of-sample performance is calculated as
@@
In out-of-sample tests, MR-DRO models consistently have the best performance among all methods, reported in Table 7.
@@
$40$ MR-DRO (Min-max) $-0.031\pm 0.491\times 10^{-3}$
MR-DRO (Exponential) $-0.034\pm 0.545\times 10^{-3}$
MR-DRO (Variable-share) $-0.034\pm 0.519\times 10^{-3}$
DRO (h1) $-0.012\pm 1.052\times 10^{-3}$
DRO (h2) $-0.012\pm 1.517\times 10^{-3}$
DRO (h3) $-0.008\pm 1.627\times 10^{-3}$
DRO (h4) $-0.025\pm 1.070\times 10^{-3}$
@@ 6 Conclusion
For future research, the entire process can be modeled, including data fusion, optimization, and trust update, within an HCI framework. In such a framework, trust can be treated as a decision variable to minimize worst-case losses more efficiently. In addition, simulation experiments could also provide deeper insight into how decision makers develop and adjust their trust. Another promising direction is to extend the framework to more complex real-world problems, such as evacuation route recommendations and emergency resource allocation, which may introduce additional computational challenges.
@@
This work is supported in part by the Army Research Laboratory under contract no. W911NF2320215.
Appendix A Proofs
@@
$\sup_{\mathbb{Q}\in\mathbb{B}_{\epsilon}(\hat{\mathbb{P}}_{HI})}\mathbb{E}^{\mathbb{Q}}\left[\ell(\bm{x},\bm{\xi})\right]$ =$=\begin{cases}\sup_{\Pi,\mathbb{Q}}\int_{\Xi}\ell(\bm{x},\bm{\xi})\mathbb{Q}(d\bm{\xi})\\ \textrm{s.t.}\quad\int_{\Xi^{2}}\|\bm{\xi}-\bm{\xi}^{{}^{\prime}}\|\Pi(d\bm{\xi},d\bm{\xi}^{{}^{\prime}})\leq\epsilon\end{cases}$ (19)
=$=\begin{cases}\sup_{\mathbb{Q}^{(i)}_{h}\in\mathcal{M}(\Xi)}\sum_{h=1}^{H}\sum_{i=1}^{I}\frac{t^{(I)}_{h}}{I}\int_{\Xi}\ell(\bm{x},\bm{\xi})\mathbb{Q}^{(i)}_{h}(d\bm{\xi})\\ \textrm{s.t.}\quad\sum_{h=1}^{H}\sum_{i=1}^{I}\frac{t^{(I)}_{h}}{I}\int_{\Xi}\|\bm{\xi}-\hat{\bm{\xi}}^{(i)}_{h}\|\mathbb{Q}^{(i)}_{h}(d\bm{\xi})\leq\epsilon,\end{cases}$ (20)
where $\Pi$ is a joint distribution of $\bm{\xi}$ and $\bm{\xi}^{{}^{\prime}}$ with marginals $\mathbb{Q}$ and ^$\hat{\mathbb{P}}_{HI}$, respectively. We drop the minimization problem in (5) and obtain (19) since the minimization of $d_{W}(\mathbb{Q},\hat{\mathbb{P}}_{HI})$ is at most radius $\epsilon$ equivalent as (19) has a feasible solution. The second equality (20) follows from the law of total probability, which means that any joint probability distribution $\Pi$ of $\bm{\xi}$ and $\bm{\xi}^{{}^{\prime}}$ can be constructed from the marginal distribution ^$\hat{\mathbb{P}}_{HI}$ of $\bm{\xi}^{{}^{\prime}}$ and the conditional distributions $\mathbb{Q}^{(i)}_{h}$ of $\bm{\xi}$ given $\bm{\xi}^{{}^{\prime}}=\hat{\bm{\xi}}^{(i)}_{h}$, for all $i\in[I]$ and $h\in[H]$. In other words, we have $\Pi=\sum_{h=1}^{H}\sum_{i=1}^{I}\frac{t^{(I)}_{h}}{I}\cdot\delta(\bm{\xi}-\hat{\bm{\xi}}^{(i)}_{h})\bigotimes\mathbb{Q}^{(i)}_{h}$, where $\bigotimes$ denotes the Kronecker product.
Applying the standard duality argument (Bertsimas and Tsitsiklis 1997), we obtain
@@
=$=\begin{cases}\inf_{\lambda,s_{hi},\bm{z}_{hij}}\quad\lambda\epsilon+\sum_{h=1}^{H}\sum_{i=1}^{I}\frac{t^{(I)}_{h}}{I}s_{hi}\\ \textrm{s.t.}\quad\sup_{\bm{\xi}\in\Xi}\left\{\ell_{j}(\bm{x},\bm{\xi})-\max_{\|\bm{z}_{hij}\|_{*}\leq\lambda}\langle\bm{z}_{hij},\bm{\xi}-\hat{\bm{\xi}}^{(i)}_{h}\rangle\right\}\leq s_{hi},\quad\forall h\in[H],i\in[I],j\in[J],\\ \quad\quad\lambda\geq 0\end{cases}$ (22b)
@@
=$=\begin{cases}\inf_{\lambda,s_{hi},\bm{z}_{hij}}\quad\lambda\epsilon+\sum_{h=1}^{H}\sum_{i=1}^{I}\frac{t^{(I)}_{h}}{I}s_{hi}\\ \textrm{s.t.}\quad[-\ell_{j}+\chi_{\Xi}]^{*}(-\bm{z}_{hij})+\langle\bm{z}_{hij},\hat{\bm{\xi}}^{(i)}_{h}\rangle\leq s_{hi},\quad\forall h\in[H],i\in[I],j\in[J],\\ \quad\quad\|\bm{z}_{hij}\|_{*}\leq\lambda,\quad\forall h\in[H],i\in[I],j\in[J],\end{cases}$ (22d)
=$=\begin{cases}\inf_{\lambda,s_{hi},\bm{z}_{hij}}\quad\lambda\epsilon+\sum_{h=1}^{H}\sum_{i=1}^{I}\frac{t^{(I)}_{h}}{I}s_{hi}\\ \textrm{s.t.}\quad[-\ell_{j}+\chi_{\Xi}]^{*}(\bm{z}_{hij})-\langle\bm{z}_{hij},\hat{\bm{\xi}}^{(i)}_{h}\rangle\leq s_{hi},\quad\forall h\in[H],i\in[I],j\in[J],\\ \quad\quad\|\bm{z}_{hij}\|_{*}\leq\lambda,\quad\forall h\in[H],i\in[I],j\in[J],\\ \end{cases}$ (22e)
We obtain (22b) from (22a) by the definition of the dual norm $\|\bm{z}_{hij}\|_{*}:=\sup_{\|\bm{u}\|\leq 1}\langle\bm{z}_{hij},\bm{u}\rangle$. This measures the maximum projection of $\bm{z}_{hij}\in\mathbb{R}^{M}$ onto any unit vector $\bm{u}\in\mathbb{R}^{M}$. By introducing the constraint $\|\bm{z}_{hij}\|_{*}\leq\lambda$, we replace $\bm{u}$ with $\bm{\xi}-\hat{\bm{\xi}}^{(i)}_{h}$ and effectively bound the influence of the deviation $\bm{\xi}-\hat{\bm{\xi}}^{(i)}_{h}$ by scaling it with $\lambda$. Then we interchange the maximization over $\bm{z}_{hij}$ with the minus sign and get an re-expressed upper bound as (22c). Finally, we reach (22d) by applying the definition of conjugacy. Recall the definition of the conjugate function. For a given $\bm{x}\in\mathbb{X}$, we have $[-\ell_{j}+\chi_{\Xi}]^{*}(\bm{v})=\sup_{\bm{\xi}\in\mathbb{R}^{M}}\left\{\langle\bm{v},\bm{\xi}\rangle+\ell_{j}(\bm{x},\bm{\xi})-\chi_{\Xi}(\bm{\xi})\right\}$. The characteristic function $\chi_{\Xi}$ is defined as $\chi_{\Xi}(\bm{\xi})=0$ if $\bm{\xi}\in\Xi$; $\chi_{\Xi}(\bm{\xi})=\infty$ otherwise. Therefore, the conjugate simplifies to $[-\ell_{j}+\chi_{\Xi}]^{*}(\bm{v})=\sup_{\bm{\xi}\in\Xi}\left\{\langle\bm{v},\bm{\xi}\rangle+\ell_{j}(\bm{x},\bm{\xi})\right\}$. Subtituting $\bm{v}=-\bm{z}_{hij}$, we have $[-\ell_{j}+\chi_{\Xi}]^{*}(-\bm{z}_{hij})=\sup_{\xi\in\Xi}\left\{\ell_{j}(\bm{x},\bm{\xi})-\langle\bm{z}_{hij},\bm{\xi}\rangle\right\}$. The equality from (22d) to (22e) holds due to a substitution of $\bm{z}_{hij}$ with $-\bm{z}_{hij}$.
Under Assumption 6, the inequality (21a) and (22c) are reduced to equalities based on Proposition 3.4 in (Shapiro 2001) and Proposition 5.5.4 in (Bertsekas 2009), leading to the equality between the optimal value of the inner worst-case expectation and (22e).
Furthermore, by Theorem 11.23(a) in Rockafellar and Wets (2009), we have
@@
=$=\begin{cases}b_{j}&\textrm{if}\quad\bm{z_{hij}}=-\bm{a}_{j},\\ \infty&\textrm{else}.\end{cases}$
@@
Proof: We first demonstrate how to reduce the inner worst-case expectation to a finite convex program under all assumptions in this context. The reformulation process up to the point where we introduce epigraphical auxiliary variables follows a similar approach to the proof of Theorem 3.1, so we have
@@
where the separability of the overall loss function enable us to interchange the summation and maximization and obtain (24b) from (24a). Then we introduce epigraphical auxiliary variables $s_{hin}$, $h\in[H]$, $i\in[I]$, $n\in[N]$:
@@
=$=\begin{cases}\inf_{\lambda,s_{hin},\bm{z}_{hijn}}\quad\lambda\epsilon+\sum_{h=1}^{H}\sum_{i=1}^{I}\sum_{n=1}^{N}\frac{t^{(I)}_{h}}{I}s_{hin}\\ \textrm{s.t.}\quad\sup_{\bm{\xi}_{n}\in\Xi_{n}}\left\{\ell_{nj}(\bm{x},\bm{\xi}_{n})-\max_{\|\bm{z}_{hijn}\|_{*}\leq\lambda}\langle\bm{z}_{hijn},\bm{\xi}_{n}-\hat{\bm{\xi}}^{(i)}_{h,n}\rangle\right\}\leq s_{hin},\quad\forall h\in[H],i\in[I],j\in[J],n\in[N]\\ \quad\quad\lambda\geq 0\end{cases}$ (25b)
$\leq\begin{cases}\inf_{\lambda,s_{hin},\bm{z}_{hijn}}\quad\lambda\epsilon+\sum_{h=1}^{H}\sum_{i=1}^{I}\sum_{n=1}^{N}\frac{t^{(I)}_{h}}{I}s_{hin}\\ \textrm{s.t.}\quad\sup_{\bm{\xi}_{n}\in\Xi_{n}}\left\{\ell_{nj}(\bm{x},\bm{\xi}_{n})-\langle\bm{z}_{hijn},\bm{\xi}_{n}\rangle\right\}+\langle\bm{z}_{hijn},\hat{\bm{\xi}}^{(i)}_{h,n}\rangle\leq s_{hin},\quad\forall h\in[H],i\in[I],j\in[J],n\in[N]\\ \quad\quad\|\bm{z}_{hijn}\|_{*}\leq\lambda,\quad\forall h\in[H],i\in[I],j\in[J],n\in[N]\\ \end{cases}$ (25c)
=$=\begin{cases}\inf_{\lambda,s_{hin},\bm{z}_{hijn}}\quad\lambda\epsilon+\sum_{h=1}^{H}\sum_{i=1}^{I}\sum_{n=1}^{N}\frac{t^{(I)}_{h}}{I}s_{hin}\\ \textrm{s.t.}\quad[-\ell_{nj}+\chi_{\Xi_{n}}]^{*}(-\bm{z}_{hijn})+\langle\bm{z}_{hijn},\hat{\bm{\xi}}^{(i)}_{h,n}\rangle\leq s_{hin},\quad\forall h\in[H],i\in[I],j\in[J],n\in[N],\\ \quad\quad\|\bm{z}_{hijn}\|_{*}\leq\lambda,\quad\forall h\in[H],i\in[I],j\in[J],n\in[N].\\ \end{cases}$ (25d)
@@
=$=\begin{cases}b_{nj}&\textrm{if}\quad\bm{z_{hijn}}=-\bm{a}_{nj},\\ \infty&\textrm{else},\end{cases}$
@@
Proof: Let $T_{i}$ be a time-inhomogeneous Markov chain representing the trust level of the source $1$ with deviation $Y$, defined on the finite state space $S=\{0,\Delta t,2\Delta t,\ldots,1\}$, $|S|=1/\Delta t+1$. Note that our assumptions about $\bm{t}^{(0)}$ and $\Delta t$ ensure that $11/\Delta t+$ 1 is a positive integer. For state $s=0$, the probability of going to state $\Delta t$ in event $i$ is $p^{\text{up}}_{0}(i)$ and the probability of staying in state 00 is $1-p^{\text{stay}}_{0}(i)$. For state $10<s<$ 1, the probability of going to state $s+\Delta t$ in event $i$ is $p^{\text{up}}_{s}(i)$, while the probability of going to state $s-\Delta t$ in event $i$ is $p^{\text{down}}_{s}(i)=1-p^{\text{up}}_{s}(i)-p^{\text{stay}}_{s}(i)$. For state $s=1$, the probability of staying in the current state is $p^{\text{stay}}_{1/\Delta t}(i)$ and the probability of going to state $s-\Delta t$ is $p^{\text{down}}_{s}(i)=1-p^{\text{stay}}_{1/\Delta t}(i)$.
@@
Since $V_{0}\leq 1$, $\Delta t>0$, and $02\beta-1>$ 0, the expected time $\mathbb{E}[\tau]$ is finite.
@@
Since the deviations $\|\Delta\xi^{(i)}_{h}\|$ are independent and bounded within $[0,L_{\text{max}}]$, we can apply Hoeffding’s inequality to bound the probability that the cumulative deviations deviates from its expected value:

### v2

@@ 5.3.2 Baseline settings for the portfolio optimization problem
We use NASDAQ100 dataset in \@BBOPcitet\@BAP\@BBNBruni et al. (2016)\@BBCP with $K_{\text{a}}=82$. The true return $\bm{r}^{(i)}_{\textrm{true}}$ at time period $i$ is a vector with each element equals the true return value corresponding to the dataset. The predicted return of asset $k$ from source $h$ at time period $i$ is sampled from a truncated normal distribution with mean $(r^{(i)}_{\textrm{true},k}+\mu_{h,k})$ and variance $\sigma^{2}_{h,k}$ within interval (1,1)(-1,1), which means no asset can lose more than $100\%$ of its value. We set $\alpha=20\%$ and $\rho=10$ in all numerical experiments. The step size for min-max error trust update is $\Delta t=0.01$. The update rate is $\eta=10^{2}$ for exponential error trust update as well as variable-share error trust update, and the share portion is $\beta=0.5$ in variable-share trust update. We conduct $30$ trials with distinct random seeds for each trust update algorithm to observe common patterns during the trust update process.
@@ 5.3.3 Baseline results for the portfolio optimization problem
MR-DRO(Min-max) $018\pm 8.223\times 10^{-6} 0.0$$002\pm 0.233\times 10^{-3}$ $472.177\pm 5.608$
MR-DRO(Exponential) $018\pm 8.225\times 10^{-6} 0.0$$002\pm 0.218\times 10^{-3}$ $444.359\pm 7.746$
MR-DRO(Variable-share) $018\pm 8.224\times 10^{-6} 0.0$$002\pm 0.219\times 10^{-3}$ $489.041\pm 6.029$
DRO (h1) $018\pm 1.324\times 10^{-5} 0.0$$004\pm 0.703\times 10^{-3}$ $247.786\pm 3.377$
DRO (h2) $018\pm 8.312\times 10^{-6} 0.0$$004\pm 0.799\times 10^{-3}$ $251.808\pm 3.679$
DRO (h3) $018\pm 1.001\times 10^{-5} 0.0$$005\pm 0.673\times 10^{-3}$ $248.473\pm 3.262$
DRO (h4) $018\pm 1.071\times 10^{-5} 0.0$$004\pm 0.635\times 10^{-3}$ $248.997\pm 3.396$
@@
We also performed out-of-sample performance analysis, where for given solution $\bm{x}^{*}$, the performance metric is given by:
@@
MR-DRO models consistently have the best performance among all methods, reported in Table 7.
@@
$40$ MR-DRO (Min-max) $0.031\pm 0.491\times 10^{-3}$
MR-DRO (Exponential) $0.034\pm 0.545\times 10^{-3}$
MR-DRO (Variable-share) $0.034\pm 0.519\times 10^{-3}$
DRO (h1) $0.012\pm 1.052\times 10^{-3}$
DRO (h2) $0.012\pm 1.517\times 10^{-3}$
DRO (h3) $0.008\pm 1.627\times 10^{-3}$
DRO (h4) $0.025\pm 1.070\times 10^{-3}$
@@ 6 Conclusion
This work is supported in part by the United States (US) Army Research Laboratory under contract no. W911NF2320215 and the US National Science Foundation (NSF) grant no. ECCS-2533775.
Appendix A Supplement Materials
@@
$\sup_{\mathbb{Q}\in\mathbb{B}_{\epsilon}(\hat{\mathbb{P}}_{HI})}\mathbb{E}^{\mathbb{Q}}\left[\ell(\bm{x},\bm{\xi})\right]$ $=\begin{cases}\sup_{\Pi,\mathbb{Q}}\int_{\Xi}\ell(\bm{x},\bm{\xi})\mathbb{Q}(d\bm{\xi})\\ \textrm{s.t.}\quad\int_{\Xi^{2}}\|\bm{\xi}-\bm{\xi}^{{}^{\prime}}\|\Pi(d\bm{\xi},d\bm{\xi}^{{}^{\prime}})\leq\epsilon\end{cases}$ (19)
$=\begin{cases}\sup_{\mathbb{Q}^{(i)}_{h}\in\mathcal{M}(\Xi)}\sum_{h=1}^{H}\sum_{i=1}^{I}\frac{t^{(I)}_{h}}{I}\int_{\Xi}\ell(\bm{x},\bm{\xi})\mathbb{Q}^{(i)}_{h}(d\bm{\xi})\\ \textrm{s.t.}\quad\sum_{h=1}^{H}\sum_{i=1}^{I}\frac{t^{(I)}_{h}}{I}\int_{\Xi}\|\bm{\xi}-\hat{\bm{\xi}}^{(i)}_{h}\|\mathbb{Q}^{(i)}_{h}(d\bm{\xi})\leq\epsilon,\end{cases}$ (20)
where $\Pi$ is a joint distribution of $\bm{\xi}$ and $\bm{\xi}^{{}^{\prime}}$ with marginals $\mathbb{Q}$ and $\hat{\mathbb{P}}_{HI}$, respectively. We drop the minimization problem in (5) and obtain (19) since the minimization of $d_{W}(\mathbb{Q},\hat{\mathbb{P}}_{HI})$ is at most radius $\epsilon$ equivalent as (19) has a feasible solution. The second equality (20) follows from the law of total probability, which means that any joint probability distribution $\Pi$ of $\bm{\xi}$ and $\bm{\xi}^{{}^{\prime}}$ can be constructed from the marginal distribution $\hat{\mathbb{P}}_{HI}$ of $\bm{\xi}^{{}^{\prime}}$ and the conditional distributions $\mathbb{Q}^{(i)}_{h}$ of $\bm{\xi}$ given $\bm{\xi}^{{}^{\prime}}=\hat{\bm{\xi}}^{(i)}_{h}$, for all $i\in[I]$ and $h\in[H]$. In other words, we have $\Pi=\sum_{h=1}^{H}\sum_{i=1}^{I}\frac{t^{(I)}_{h}}{I}\cdot\delta(\bm{\xi}-\hat{\bm{\xi}}^{(i)}_{h})\bigotimes\mathbb{Q}^{(i)}_{h}$, where $\bigotimes$ denotes the Kronecker product.
Applying the standard duality argument \@BBOPcitep\@BAP\@BBN(Bertsimas and Tsitsiklis 1997)\@BBCP, we obtain
@@
$=\begin{cases}\inf_{\lambda,s_{hi},\bm{z}_{hij}}\quad\lambda\epsilon+\sum_{h=1}^{H}\sum_{i=1}^{I}\frac{t^{(I)}_{h}}{I}s_{hi}\\ \textrm{s.t.}\quad\sup_{\bm{\xi}\in\Xi}\left\{\ell_{j}(\bm{x},\bm{\xi})-\max_{\|\bm{z}_{hij}\|_{*}\leq\lambda}\langle\bm{z}_{hij},\bm{\xi}-\hat{\bm{\xi}}^{(i)}_{h}\rangle\right\}\leq s_{hi},\quad\forall h\in[H],i\in[I],j\in[J],\\ \quad\quad\lambda\geq 0\end{cases}$ (22b)
@@
$=\begin{cases}\inf_{\lambda,s_{hi},\bm{z}_{hij}}\quad\lambda\epsilon+\sum_{h=1}^{H}\sum_{i=1}^{I}\frac{t^{(I)}_{h}}{I}s_{hi}\\ \textrm{s.t.}\quad[-\ell_{j}+\chi_{\Xi}]^{*}(-\bm{z}_{hij})+\langle\bm{z}_{hij},\hat{\bm{\xi}}^{(i)}_{h}\rangle\leq s_{hi},\quad\forall h\in[H],i\in[I],j\in[J],\\ \quad\quad\|\bm{z}_{hij}\|_{*}\leq\lambda,\quad\forall h\in[H],i\in[I],j\in[J],\end{cases}$ (22d)
[…]
We obtain (22b) from (22a) by the definition of the dual norm $\|\bm{z}_{hij}\|_{*}:=\sup_{\|\bm{u}\|\leq 1}\langle\bm{z}_{hij},\bm{u}\rangle$. This measures the maximum projection of $\bm{z}_{hij}\in\mathbb{R}^{M}$ onto any unit vector $\bm{u}\in\mathbb{R}^{M}$. By introducing the constraint $\|\bm{z}_{hij}\|_{*}\leq\lambda$, we replace $\bm{u}$ with $\bm{\xi}-\hat{\bm{\xi}}^{(i)}_{h}$ and effectively bound the influence of the deviation $\bm{\xi}-\hat{\bm{\xi}}^{(i)}_{h}$ by scaling it with $\lambda$. Then we interchange the maximization over $\bm{z}_{hij}$ with the minus sign and get an re-expressed upper bound as (22c). Finally, we reach (22d) by applying the definition of conjugacy. Recall the definition of the conjugate function. For a given $\bm{x}\in\mathbb{X}$, we have $\ell_{j}+\chi_{\Xi}]^{*}(\bm{v})=\sup_{\bm{\xi}\in\mathbb{R}^{M}}\left\{\langle\bm{v},\bm{\xi}\rangle+\ell_{j}(\bm{x},\bm{\xi})-\chi_{\Xi}(\bm{\xi})\right\}$. The characteristic function $\chi_{\Xi}$ is defined as $\chi_{\Xi}(\bm{\xi})=0$ if $\bm{\xi}\in\Xi$; $\chi_{\Xi}(\bm{\xi})=\infty$ otherwise. Therefore, the conjugate simplifies to $\ell_{j}+\chi_{\Xi}]^{*}(\bm{v})=\sup_{\bm{\xi}\in\Xi}\left\{\langle\bm{v},\bm{\xi}\rangle+\ell_{j}(\bm{x},\bm{\xi})\right\}$. Subtituting $\bm{v}=-\bm{z}_{hij}$, we have $\ell_{j}+\chi_{\Xi}]^{*}(-\bm{z}_{hij})=\sup_{\xi\in\Xi}\left\{\ell_{j}(\bm{x},\bm{\xi})-\langle\bm{z}_{hij},\bm{\xi}\rangle\right\}$. The equality from (22d) to (22e) holds due to a substitution of $\bm{z}_{hij}$ with $\bm{z}_{hij}$.
Under Assumption 6, the inequality (21a) and (22c) are reduced to equalities based on Proposition 3.4 in \@BBOPcitep\@BAP\@BBN(Shapiro 2001)\@BBCP and Proposition 5.5.4 in \@BBOPcitep\@BAP\@BBN(Bertsekas 2009)\@BBCP, leading to the equality between the optimal value of the inner worst-case expectation and (22e).
Furthermore, by Theorem 11.23(a) in \@BBOPcitet\@BAP\@BBNRockafellar and Wets (2009)\@BBCP, we have
@@
$=\begin{cases}b_{j}&\textrm{if}\quad\bm{z_{hij}}=-\bm{a}_{j},\\ \infty&\textrm{else}.\end{cases}$
@@
Proof: We first demonstrate how to reduce the inner worst-case expectation to a finite convex program under all assumptions in this context. The reformulation process up to the point where we introduce epigraphical auxiliary variables follows a similar approach to the proof of Theorem 3.1, and thus,
@@
where the separability of the overall loss function enable us to interchange the summation and maximization and obtain (24b) from (24a). We then introduce epigraphical auxiliary variables $s_{hin}$, $h\in[H]$, $i\in[I]$, $n\in[N]$:
@@
$=\begin{cases}\inf_{\lambda,s_{hin},\bm{z}_{hijn}}\quad\lambda\epsilon+\sum_{h=1}^{H}\sum_{i=1}^{I}\sum_{n=1}^{N}\frac{t^{(I)}_{h}}{I}s_{hin}\\ \textrm{s.t.}\quad\sup_{\bm{\xi}_{n}\in\Xi_{n}}\left\{\ell_{nj}(\bm{x},\bm{\xi}_{n})-\max_{\|\bm{z}_{hijn}\|_{*}\leq\lambda}\langle\bm{z}_{hijn},\bm{\xi}_{n}-\hat{\bm{\xi}}^{(i)}_{h,n}\rangle\right\}\leq s_{hin},\quad\forall h\in[H],i\in[I],j\in[J],n\in[N]\\ \quad\quad\lambda\geq 0\end{cases}$ (25b)
[…]
@@
$=\begin{cases}b_{nj}&\textrm{if}\quad\bm{z_{hijn}}=-\bm{a}_{nj},\\ \infty&\textrm{else},\end{cases}$
@@
Proof: Let $T_{i}$ be a time-inhomogeneous Markov chain representing the trust level of the source $1$ with deviation $Y$, defined on the finite state space $S=\{0,\Delta t,2\Delta t,\ldots,1\}$, $|S|=1/\Delta t+1$. Note that our assumptions about $\bm{t}^{(0)}$ and $\Delta t$ ensure that $1/\Delta t+1$ is a positive integer. For state $s=0$, the probability of going to state $\Delta t$ in event $i$ is $p^{\text{up}}_{0}(i)$ and the probability of staying in state 0 is $1-p^{\text{stay}}_{0}(i)$. For state $0<s<1$, the probability of going to state $s+\Delta t$ in event $i$ is $p^{\text{up}}_{s}(i)$, while the probability of going to state $s-\Delta t$ in event $i$ is $p^{\text{down}}_{s}(i)=1-p^{\text{up}}_{s}(i)-p^{\text{stay}}_{s}(i)$. For state $s=1$, the probability of staying in the current state is $p^{\text{stay}}_{1/\Delta t}(i)$ and the probability of going to state $s-\Delta t$ is $p^{\text{down}}_{s}(i)=1-p^{\text{stay}}_{1/\Delta t}(i)$.
@@
Since $V_{0}\leq 1$, $\Delta t>0$, and $2\beta-1>0$, the expected time $\mathbb{E}[\tau]$ is finite.
@@
Since the deviations $\|\Delta\xi^{(i)}_{h}\|$ are independent and bounded within $0,L_{\text{max}}]$, we can apply Hoeffding’s inequality to bound the probability that the cumulative deviations deviates from its expected value: